OpenStudy (anonymous):

Calculate the following integral \[\int\limits x^3 * e^x^2\]

OpenStudy (mimi_x3):

\[\int x^{3}*e^{x^2}dx\]

OpenStudy (anonymous):

Yes that's it, don't know why it wasn't showing that

OpenStudy (mimi_x3):

\[ u = x^2\] du / dx = 2x

OpenStudy (mimi_x3):

\[ \frac{1}{2} \int ue^{u}du\]

OpenStudy (mimi_x3):

integration by parts.

OpenStudy (mimi_x3):

i think you can do it now?

OpenStudy (anonymous):

If u = x^3 then isn't du/dx= 3x^2

OpenStudy (mimi_x3):

no, cos e^{x^2}

OpenStudy (mimi_x3):

but you can do u = x^2, if you want.